OpenStudy (anonymous):

sin30 degree cos30 degree +cos30 degree sin30 degree which expression do i use in here and wht is the exact answer ?

OpenStudy (anonymous):

\[\sin(\theta_{1}+\theta_{2)}=\sin(\theta_{1})\cos(\theta_{2})+\cos(\theta_{1})\sin(\theta_{2})\]

OpenStudy (anonymous):

sin30 degree cos30 degree +cos30 degree sin30=sin(30+30) =sin(60 degree)= \[\frac{ \sqrt{3} }{ 2 }\]

OpenStudy (usukidoll):

triangles